SourMunchkin7806:

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = x3 + 4 and g(x) = Cube root of quantity x minus four.

SourMunchkin7806:

@Arcadiouse

Arcadiouse:

Cant help you

SourMunchkin7806:

know anyone who can?

dude:

\(\large {f(x) = x^3 + 4 ~and~ g(x) = \sqrt[3]{x-4}}\) \(f(g(x))=\boxed{\sqrt[3]{x-4}}^3+4=\color{red}?\) and \(g(f(x))=\sqrt[3]{\boxed{x^3+4}-4}=\color{red}?\)